I really need help with all 3 please, I will give a good rating.

1. A 5.0 gram nickel has been placed on a large rotating disk a distance 12 centimeters from the center. The disk spins at a rate of 33 revolutions per minute. How long will it take the nickel to rotate through 8.6 radians?

2. A 1200 kg car is traveling around a circular track of diameter 0.48 kilometers. If the car travels an angle of 39 degrees around the track in 6.0 seconds, calculate the tangential speed of the car.

3. A 58 kg person is standing on a scale in the bed of a pickup truck. At the bottom of a large circular valley in the road, where radius of curvature is 96 meters, the truck is moving forward with a speed of 110 km/hour. What does the scale read, in Newtons?
Answer #1

1.

angular velocity

w = 33* 2* 3.14 / 60 = 3.454 rad/s

now as we know

t = d/w = 8.6 /3.454 = 2.49 s

====

2.

distance traveld by car

d = ( x/360) (2pi r)

d = (39/360)* 2* 3.14* 0.24* 1000

d = 163.28 m

now velocity

v = d/t = 163.28 / 6 = 27.213 m/s

=====

3.

scale reading

W = mg + m v^2/r

W = 58* ( 9.8 + (110* 5/18)^2 / 96)

W = 1132.475 N

====

Comment before rate in case any doubt, will reply for sure.. goodluck